OpenStudy (anonymous):

I dont know this one... Tesha can type 30 words per minute how long will it take her to type her 3-page paper if each page is about 250 words long

OpenStudy (saifoo.khan):

total words = 3 * 250 = 750

OpenStudy (saifoo.khan):

now, 750/ 30 75/3 25mins

OpenStudy (anonymous):

so 750? :D

OpenStudy (saifoo.khan):

No, 25mins. 750 are the total number of words.

OpenStudy (anonymous):

oooh

OpenStudy (anonymous):

First Saif found how many words Tesha will need to type. There are 3 pages and each page has 250 words. So together all 3 pages contain: 3 (pages) * 250 (words on a page) = 750 words So she will need to type 750 words. Is this part clear ?

OpenStudy (anonymous):

750 words divided by 30 words/min =25 minutes